OpenStudy (anonymous):

The Mendes family bought a new house 11 years ago for $100,000. The house is now worth $196,000. Assuming a steady rate of growth, what was the yearly rate of appreciation? Round your answer to the nearest tenth of a percent (1.2% etc)

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

growth : y = a(1+r)^2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[196000 = 100000 (1 + \frac{ r }{ 100 } ) ^{11}\]

OpenStudy (anonymous):

would i multiplu 100 to both sides ?

OpenStudy (anonymous):

divide by 100000 on both sides

OpenStudy (anonymous):

1.96

OpenStudy (phi):

you probably should be using the continuous interest formula http://cs.selu.edu/~rbyrd/math/continuous/ \[ P= P_0 e^{r \ t} \]

OpenStudy (anonymous):

\[1.96 = (1+\frac{ r }{ 100 })^{11}\]

OpenStudy (anonymous):

now can i divide by 100

OpenStudy (anonymous):

no you need to take log

OpenStudy (phi):

if you assume compounding at a yearly rate, your answer will be off by a few tenths of a percent

OpenStudy (anonymous):

yea it asked for yearly rate of appreciation

OpenStudy (anonymous):

log1.96=log (1+r/100)^2

OpenStudy (anonymous):

^11****

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how can there be power after you log

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[\log_{11} (1 + \frac{ r }{ 100 }) = 1.96 \]

OpenStudy (anonymous):

\[\log_{11} (1.96) = 0.28064\]

OpenStudy (phi):

***Assuming a steady rate of growth*** means continuous they want the annual interest rate assuming continuous compounding.

OpenStudy (anonymous):

yea thats what we've been using

OpenStudy (anonymous):

okay i about to solve it tell me if im correct

OpenStudy (anonymous):

2.8%?

OpenStudy (anonymous):

I believe you are to assume annual compounding. By "steady rate of growth" I believe the author simply means that the annual rate of appreciation is constant over the duration of the problem.

OpenStudy (anonymous):

no not 2.8

OpenStudy (anonymous):

dayum if thats the case its simple interest formula

OpenStudy (anonymous):

you move the decimal two places forward or back

OpenStudy (anonymous):

And base 11 logs are not required. You have\[1.96=\left( 1+\frac{ r }{ 100 } \right)^{11}\]Using base 10 logs\[\log1.96 = 11 \log \left( 1+\frac{ r }{ 100 } \right)\]\[\frac{ \log 1.96 }{ 11 } = \log \left( 1+\frac{ r }{ 100 } \right)\]\[0.02657 = \log \left( 1+\frac{ r }{ 100 } \right)\]\[1+\frac{ r }{ 100 } = 10^{0.02657} = 1.0631\]Can you take it from here and solve for \(r\)?

OpenStudy (anonymous):

no i dont understand

OpenStudy (anonymous):

would it be .0631/100 ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

0.000631 = r

OpenStudy (anonymous):

So close. \[1+\frac{ r }{ 100 } = 0.0631\]Subtract 1 from both sides\[\frac{ r }{ 100 } = 0.0631\]How do you solve for r?

OpenStudy (anonymous):

6.31

OpenStudy (anonymous):

Sorry. First line should read\[1+\frac{ r }{ 100 } = 1.0631\]

OpenStudy (anonymous):

6.31%

OpenStudy (anonymous):

You are correct. Now round to the number of decimal places required. Well done?

OpenStudy (anonymous):

Well done!

OpenStudy (anonymous):

now you need to round off somemore

OpenStudy (anonymous):

can yall help me with another ?

OpenStudy (anonymous):

Delete this question and post the other.

OpenStudy (anonymous):

it said that was wrong . it says to round to the nearest tenth of percent

OpenStudy (anonymous):

6.3% correct ??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i told you to round off somemore lol

OpenStudy (anonymous):

it still says its wrong

OpenStudy (anonymous):

We both indicated that you were to round to the correct number of decimal places :)

OpenStudy (anonymous):

its still wrong . i put 6.3%

OpenStudy (anonymous):

Well. I don't know what the issue is. \[196000 = 100000\times 1.603^{11}\]so the math checks out.

OpenStudy (anonymous):

Sorry. 1.063^11

OpenStudy (anonymous):

wanna try simple interest method?

OpenStudy (anonymous):

I can't imagine that it's continuous compounding but, if so, you'll have to use natural logs.\[F=Pe^{rt}\]\[196000 = 100000e^{11r}\]\[1.96 = e^{11r}\]\[\ln 1.96 = 11r\]\[r=\frac{ \ln 1.96 }{ 11 }\]

OpenStudy (anonymous):

okay .

OpenStudy (anonymous):

.06117

OpenStudy (anonymous):

Convert to percent by multiplying by 100% and round to the correct number of decimal places.

OpenStudy (anonymous):

hate that kinda english-testing maths questions

OpenStudy (phi):

yes, you get 6.1%

OpenStudy (anonymous):

Me too. Just finished a graduate course in financial mathematics and "steady rate of growth" was never used to indicate continuous compounding. @phi , if this is the correct answer, I owe you an apology. Please forgive my doubt.

OpenStudy (anonymous):

it was correct

OpenStudy (anonymous):

Congratualtions

OpenStudy (anonymous):

sorta can't understand how e comes into play in continuous compounding

OpenStudy (anonymous):

gratz

OpenStudy (anonymous):

THANK YALL !!!! . can yall help me with the next one